write an equation in standard form for the line that is parallel to the graph of -8x= 5-4y and has a y-intercept of -0.5
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

-8x= 5-4y and has a y-intercept of -0.5
re write in slope intercept form
4y=8x+5
/4
y=2x+(5/4)
slope = 2
a parallel line will have same slope = 2
y intercept = -0.5
equation
y=2x-0.5
y=2x-1/2
multiply by 2
2y=4x-1
4x-2y=1 is the equation

Answer by stanbon(75887)

write an equation in standard form for the line that is parallel to the graph of -8x= 5-4y and has a y-intercept of -0.5
-------------
Solve the given equation for "y":
y = 2x - (5/4)
slope = 2
----------
Find equation with slope = 2 that passes thru (0,-0.5)
----
Equation: y = 2x - (5/4)
----
Change to standard form:
4y = 8x - 5
------
8x - 4y = 5
================
Cheers,
Stan H.
=================